### v1

# Self-supervised pre-training with diffusion model for few-shot landmark detection in x-ray images

Abstract
In the last few years, deep neural networks have been extensively applied in the medical domain for different tasks, ranging from image classification and segmentation to landmark detection. However, the application of these technologies in the medical domain is often hindered by data scarcity, both in terms of available annotations and images. This study introduces a new self-supervised pre-training protocol based on diffusion models for landmark detection in x-ray images. Our results show that the proposed self-supervised framework can provide accurate landmark detection with a minimal number of available annotated training images (up to $50$), outperforming ImageNet supervised pre-training and state-of-the-art self-supervised pre-trainings for three popular x-ray benchmark datasets. To our knowledge, this is the first exploration of diffusion models for self-supervised learning in landmark detection, which may offer a valuable pre-training approach in few-shot regimes, for mitigating data scarcity.
1 Introduction
Landmark detection is the task of identifying anatomical keypoints of interest in images [1]. Detecting key points can directly assist in angle measurements [2], skeletal measurements [3], and surgical planning [4].
In recent years, deep neural networks have become extensively utilized for the landmark detection task [5, 6], and many fully supervised methods have been proposed [7, 8, 9, 10, 11, 12, 13, 14, 15]. However, the limited availability of annotated data and the reliance on medical experts for annotations present challenges in applying deep learning to this domain. Specifically, annotating landmarks is labor-intensive, costly, and demands expert precision [8, 7, 9], particularly for critical tasks such as surgical treatment planning. Consequently, real-world datasets typically contain very few annotated images, requiring the design of label-efficient training solutions [16].
A transfer learning framework, typically fine-tuning ImageNet pre-trained models on medical downstream tasks, has been a common strategy to tackle this challenge [17, 18, 19]. Whether specific in-domain pre-training may benefit the model performance on the downstream task is still debated, and recent works provide controversial results depending on the specific task and domain [20, 9]. Nevertheless, involving in-domain data may be an effective alternative to ImageNet pre-training, especially in a self-supervised fashion to overcome the limited availability of annotations.
Recent advances in self-supervised learning (SSL) methods such as MoCoV3 [21], SimCLRV2 [22], and DINO [23] have shown significant potential. These approaches enable models to learn robust representations from unlabeled data, thereby reducing the dependency on large labeled training datasets. In this context, our main contribution is the proposal of a novel few-shot self-supervised pre-training approach, based on denoising diffusion probabilistic models (DDPMs), specifically focusing on landmark detection for x-ray images, widely used for anatomical assessment and diagnosis, in a realistic and typical scenario where the number of available annotations and images is minimal (up to $50$).
To our knowledge, this work is the first to employ diffusion models for self-supervision in landmark detection. We compare the proposed method with state-of-the-art alternatives including ImageNet supervised pre-training, MoCoV3, SimCLRV2, and DINO self-supervised pre-training, in terms of performance and robustness at the varying of labeled training available samples. Our results show that our method outperforms ImageNet supervised pre-training and the other self-supervised methods for the landmark detection task. Furthermore, we perform a set of experiments to provide insights on whether the usage of a different in-domain pre-training dataset may benefit the downstream task, in a few shot regimes.
The remainder of the paper is organized as follows: Section 2 reviews related studies, particularly focusing on self-supervised pre-training in the medical domain and landmark detection. Section 3 describes our proposed approach. Section 4 gives an overview of the benchmark datasets employed in this research, details the implementation specifics, and explains the evaluation metrics. Lastly, Section 5 showcases our experimental results and insights, drawing some conclusions in Section 6.
2 Related Works
The problem of few-shot learning is particularly relevant in the medical application domain, as obtaining both images and high-quality, unbiased annotations is costly and time-consuming. Consequently, several works leverage self-supervised pre-training for label efficiency [24, 25, 26, 27].
Zhu et al.[27] propose UOD (Universal One-shot Detection), a domain-adaptive framework for one-shot anatomical landmark detection across multiple medical imaging domains. UOD uses contrastive learning and a domain-adaptive transformer, while reducing annotation burden. Rousseau et al. [24] introduce a pre-training method using diffusion models for dental radiography segmentation. Baranchuk et al. [25] use diffusion models to capture high-level semantic information for semantic segmentation. Brempong et al. [26] propose Decoder Denoising Pretraining (DDeP) to enhance label efficiency in semantic segmentation models by emphasizing the importance of pre-training both encoder and decoder components with limited labeled examples.
In the specific context of landmark detection, the existing works have mainly tackled the data scarcity issue by leveraging transfer learning techniques [28, 8, 9]. Tiulpin et al. [28] use transfer learning from low-budget annotations for knee x-ray landmark localization in osteoarthritis stages. Di Via et al. [9] study whether in-domain data provide benefits over ImageNet pre-training for landmark detection in x-ray images. Zhu et al. [8] propose GU2Net, a universal model for anatomical landmark detection in medical images, addressing limitations in specialized, dataset-dependent methods. Unlike the previously cited works, in this paper, we examine the impact of self-supervised pre-training with diffusion models on landmark detection for x-ray images, focusing on a realistic and typical scenario where the number of available annotations and images is limited. To our knowledge, this is the first application of diffusion models for SSL in this specific task.
Figure 1: Schematic representation of our DDPM self-supervised landmark detection pipeline.
3 Approach
Fig. 1 illustrates our two-step methodology: initially, a DDPM is pre-trained on the training set, in a self-supervised fashion (with no annotations). Subsequently, this model is fine-tuned on the labeled training set for the downstream task, represented by the landmark detection. Further details on these steps are discussed in the following section.
Pre-training using the DDPM: Denoising Diffusion Probabilistic Models [29] (DDPMs) have emerged as a powerful class of generative models, excelling in the synthesis of high-fidelity samples in diverse fields, including medical imaging [30, 31]. Their training methodology involves a forward diffusion process that gradually transforms data into noise and a reverse process that learns to reverse this transformation, thereby generating new data samples from noise.
In this work, we leverage a DDPM Unet architecture as a predictive model for noise. During pre-training, the DDPM Unet receives batches of images alongside their respective diffusion time steps, and it is tasked with predicting the perturbation in the data. Real-world landmark detection datasets are typically small-scale, thus, we expect our DPPM Unet to be pre-trained with small-scale unannotated datasets, while still being able to provide rich and general features for the downstream task. This allows us to keep computational resources and time under a budget, while providing an effective few-shot learning approach for landmark detection.
Fine-tuning on the downstream task: The pre-trained Unet model integrated into the DDPM undergoes fine-tuning for landmark detection in a few-shot regime, thus using a few available annotated data for the downstream task. This step requires only modifying the final classification layer to align with the number of landmarks to be predicted. Although the first stage adopts a self-supervised strategy (with no annotations), the second phase employs a supervised approach by providing the model with ground-truth heat maps for each landmark to be predicted.
4 Experimental Setup
In this section, we describe the datasets used, the metrics employed, and the details of the implementation of our DDPM architectures and training procedures for the different pre-training methods discussed above. Afterwards, we show the results that have been achieved by evaluating the approach presented.
4.1 Datasets Overview
This section outlines the datasets employed in our study, including Chest x-rays, Cephalometric x-rays, and Hand x-rays. We describe the dataset composition, annotation protocols, and the division into training, validation, and testing sets to ensure clarity and reproducibility in our experiments.
Chest dataset: This dataset comprises $279$ Chest x-ray images sourced from the China set, as documented in [32], and is accessible through the U.S. National Library of Medicine’s public repository. The original image dimensions are approximately $3000$x3000300030003000 pixels. Following the procedures provided in [8, 9], we split the dataset with the first $195$ images for training, the subsequent $34$ for validation and the final $50$ for testing. We evaluate our model’s performance by pixel distance using $6$ manually annotated landmarks per image as all the PNG dataset has no information on physical spacing.
Cephalometric dataset: The Cephalometric dataset [33] is an open-source resource containing $400$ cephalometric x-ray images of individuals aged between $7$ and $76$ years. Each image has a resolution of $0.1$mmx0.10.10.10.1mm and dimensions of $2400$x1935193519351935 pixels. In our research, the first $150$ images are used for training purposes, while the remaining $250$ images are set aside for testing, following the methodology in [8]. To ensure accurate anatomical landmark annotations, two orthodontists identified $19$ landmarks on each image, and the mean coordinates were calculated to address inter-observer variability.
Hand dataset: This open dataset entails $909$ Hand x-ray images, each at an average size of $1563$x2169216921692169 pixels. Consistent with the approaches taken by [8, 9], we allocate the first $550$ images for training, the subsequent $59$ for validation, and the final $300$ images for testing. To enable a direct comparison with other approaches in terms of physical distances we take the estimate made by Payer et al.[34] that the width of the wrist is $50$mm. Then, physical distance is computed as $\frac{50}{|p-q|_{2}}$, where $p$ and $q$ are the two ends of the wrist, determined by first and fifth landmarks in the set of $37$ manually labelled by Payer et al. [34] as well.
(a) Chest landmarks: 6
(b) Cephalometric landmarks: 19
(c) Hand landmarks: 37
Figure 2: Annotated landmarks in x-ray images for Chest, Cephalometric, and Hand datasets.
4.2 Implementation Details
This section provides the technical specifics of the implementation and experimental setup. The experiments were conducted on a single NVIDIA A30 GPU with 24 GB of RAM, using Python 3.10 and PyTorch 2.1.0.
Data pre-processing: Images are resized to $256$x256256256256 pixels and normalized to a $[0-1]$ range. Post-resizing, random augmentations are applied during training, including rotations (uniformly between $-2$°and $2$°), scaling (from $-0.02$ to $0.02$), and translations (from $-0.02$ to $0.02$).
Pre-training stage: The DDPM is trained with a batch size of $4$, with gradient accumulation every $8$ batches. Training involves $500$ diffusion steps with a linear noise schedule. The training process spans $10k$ iterations, and we save the weights at iterations $4k$, $6k$, $8k$, and $10k$ for comparison during fine-tuning. We utilize the AdamW optimizer with a learning rate of $1e^{-4}$ and apply the exponential moving average to model parameters with a decay factor of $0.995$. We employ mean squared error as the loss function, consistent with the paper [29].
The DDPM Unet architecture features a multi-scale approach with channel multipliers [$1$, $2$, $4$, $8$], indicating the increase in channels at each down-sampling stage. It incorporates attention mechanisms at a resolution of $32$, uses $4$ attention heads per channel, and includes $4$ residual blocks at each resolution level.
Concerning the implementation of other SSL and ImageNet pre-training techniques, the sources used are [35] and [36].
Fine-tuning stage: The fine-tuning process is carried out using the AdamW optimizer for $200$ epochs, with an early stopping criterion. The learning rate is adjusted using a ReduceLROnPlateau scheduler based on the validation loss, starting at an initial value of $1e^{-5}$. The batch size is set at $2$, with a gradient accumulation of $8$. We use cross-entropy loss for our semantic landmark detection task to predict different classes.
4.3 Evaluation Metrics
In our research, we adopt the same evaluation metrics chosen by [8, 28, 9, 7], namely Mean Radial Error (MRE) and Success Detection Rate (SDR), which are widely used to assess the performance of landmark detection algorithms.
The MRE measures the accuracy of predicted landmarks by calculating the average Euclidean distance to the corresponding ground truth landmarks. For $n$ landmarks, the MRE is defined as:
$MRE=\frac{1}{n}\sum_{i=1}^{n}\sqrt{(x_{i}-\hat{x}_{i})^{2}+(y_{i}-\hat{y}_{i})^{2}}$
where $(x_{i},y_{i})$ represent the ground truth landmark coordinates and $(\hat{x}_{i},\hat{y}_{i})$ represent the predicted coordinates.
SDR, in contrast, evaluates robustness by indicating the percentage of predicted landmarks that fall within a certain threshold distance from the actual landmarks. For $n$ landmarks with a threshold $t$, the SDR is calculated as:
$SDR=\frac{1}{n}\sum_{i=1}^{n}I(d_{i}\leq t)$
where $d_{i}$ is the Euclidean distance between the ground truth and the predicted landmark, and $I$ is an indicator function that returns $1$ if $d_{i}\leq t$ and 00 otherwise.
5 Results and Analysis Discussion
In this section, we present the experiments conducted to evaluate our proposed methods and the corresponding results. We detail the tuning of DDPM pre-training iterations, assess the performance for landmark detection with few labeled data, and analyze the impact of different pre-training datasets on the downstream task.
5.1 Tuning the DDPM pre-training iterations
The number of training iterations during the pre-training of the DDPM is a fundamental hyperparameter of our proposed approach. To evaluate the robustness and select the number of training iterations, we perform an experiment running three times the entire pipeline, considering $4k$, $6k$, $8k$, and $10k$ training iterations for the DDPM pre-training step.
Figure 3 visualizes the landmark detection results obtained in the validation sets for the Chest, Cephalometric, and Hand datasets, in terms of MRE. Our findings suggest that the pre-training is quite robust, except for the $4k$ iterations where the model is likely still learning crucial data features (as shown by the high standard deviation). Excessive training iterations beyond $8k$ may result in overfitting, and a slightly degrading performance on the downstream task. Considering these results, we decided to select the optimal number of training iterations for DDPM pre-training for subsequent experiments, utilizing $6k$ iterations for the Chest dataset, $8k$ iterations for the Cephalometric dataset, and $8k$ iterations for the Hand dataset.
Figure 3: Landmark detection performance on Chest, Cephalometric, and Hand validation sets across DDPM pre-training iterations.
5.2 Downstream task performance evaluation
In this section, we assess the effectiveness of our DDPM self-supervised pre-training method, benchmarking it with supervised ImageNet pre-training, and self-supervised state-of-the-art methods including MoCoV3, SimCLRV2, and DINO, across different numbers of labeled training samples (up to $50$) in the Chest, the Cephalometric, and the Hand datasets. Tables 1, 2, and 3 report the average test results for three independent runs on the three datasets (a bar chart version of the tables is provided in the Supplementary Material). With the exception of our approach, which implements a DDPM as described in the original paper [29], all reported results utilize DenseNet161 as the encoder backbone. We opt for this backbone due to its performance in the landmark detection task across all the datasets tested, based on a hold-out validation comparison with VGG19 and ResNeXt50_32x4D. Our approach outperforms both ImageNet and the explored alternative SSL approaches for all the tested datasets and quantities of training images in the downstream task. These results are qualitatively supported by Figure 4, providing a visual representation (i.e., the prediction heatmaps) of the methods’ performance with $10labels$, for each of the datasets included in our work.
Performance gains are particularly pronounced in low-data regimes, crucial for medical imaging applications where labeled data is often scarce.
In the Chest dataset (Table 1), with just one labeled sample, our method achieves an MRE of $14.99$ px compared to $143.67$ px for ImageNet and $64.69$ px for the best alternative (DINO). This represents a $89.6$% reduction in MRE versus ImageNet and a $76.8$% reduction versus DINO. The SDR at $9$px shows even more improvements, with our method reaching $64.63$%, far surpassing ImageNet (00%) and DINO ($23.44$%).
Similar trends are observed in the Cephalometric dataset (Table 2). Using one labeled sample, our method attains an MRE of $6.87$ mm versus $48.69$ mm for ImageNet and $24.39$ mm for the best alternative (MoCoV3) - reductions of $85.9$% and $71.8$%, respectively. The SDR at the restrictive 2mm further highlights our method’s superiority, achieving $49.43$% compared to ImageNet’s $0.12$% and MoCoV3’s $18.36$%.
The Hand dataset (Table 3) results reinforce these findings. With one labeled sample, our approach yields an MRE of $28.75$ mm versus $79.32$ mm for ImageNet and $97.31$ mm for the best alternative (DINO) - reductions of $63.8$% and $70.5$%, respectively. The SDR at 2mm shows our method at $27.87$%, significantly outperforming both ImageNet ($0.04$%) and DINO ($0.47$%).
As the number of labeled samples increases beyond $10$, the performance gap narrows, though our approach still maintains an edge. For instance, in the Cephalometric dataset with $25$ labeled samples, our method’s MRE of $1.62$ mm is closer to, but still better than, ImageNet’s $4.18$ mm and SimCLRV2’s $2.19$ mm. This trend is consistent across all datasets, indicating that while our method provides substantial benefits in extremely low-data scenarios, for which it is conceived and proposed, it continues to offer improvements even as more labeled data becomes available.
The obtained results highlight the efficacy of our DDPM self-supervised pre-training approach for landmark detection in few-shot learning scenarios, prevalent in medical imaging contexts where available high-quality annotated images are typically limited.
Finally, to frame our results in the state-of-the-art, we can consider that the current best-performing method [9] for the landmark detection task of the whole Chest, Cephalometric, and Hand datasets, consists of fine-tuning an ImageNet pre-trained model on the target datasets.
Table 1: Comparing landmark detection test performance on the Chest dataset through our DDPM self-supervised pre-training and ImageNet, MoCoV3, SimCLRV2, DINO pre-training across varying numbers of labeled training images, highlighting best results in bold.
Number of Labeled Training Samples Pre-trained Models Chest
MRE $\downarrow$ SDR(%) $\uparrow$
(px) 3px 6px 9px
$1$ ImageNet $143.67\pm 0.27$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
MoCoV3 $106.74\pm 14.95$ $2.98\pm 0.51$ $12.33\pm 2.26$ $19.92\pm 3.47$
SimCLRV2 $118.97\pm 1.84$ $1.90\pm 0.19$ $6.09\pm 0.58$ $11.39\pm 0.58$
DINO $64.69\pm 0.91$ $3.66\pm 0.33$ $15.72\pm 0.38$ $23.44\pm 0.38$
Our method $\mathbf{14.99\pm 0.00}$ $\mathbf{19.92\pm 0.00}$ $\mathbf{46.34\pm 0.00}$ $\mathbf{64.63\pm 0.00}$
$5$ ImageNet $72.81\pm 2.71$ $2.44\pm 0.33$ $8.54\pm 1.00$ $14.09\pm 1.38$
MoCoV3 $18.35\pm 0.18$ $14.63\pm 1.00$ $41.60\pm 0.69$ $66.53\pm 0.69$
SimCLRV2 $30.09\pm 0.35$ $15.72\pm 0.69$ $48.37\pm 1.45$ $67.48\pm 0.33$
DINO $24.48\pm 1.73$ $11.25\pm 1.16$ $32.65\pm 0.51$ $48.78\pm 1.00$
Our method $\mathbf{6.04\pm 0.00}$ $\mathbf{32.52\pm 0.00}$ $\mathbf{65.04\pm 0.00}$ $\mathbf{86.99\pm 0.00}$
$10$ ImageNet $35.83\pm 1.57$ $7.11\pm 0.84$ $22.56\pm 1.06$ $37.30\pm 1.39$
MoCoV3 $13.52\pm 0.67$ $18.49\pm 0.21$ $44.72\pm 1.21$ $67.28\pm 1.43$
SimCLRV2 $28.50\pm 0.00$ $17.89\pm 0.00$ $45.93\pm 0.00$ $64.63\pm 0.00$
DINO $14.96\pm 0.03$ $16.26\pm 0.41$ $39.64\pm 0.21$ $61.99\pm 0.61$
Our method $\mathbf{5.33\pm 0.00}$ $\mathbf{33.74\pm 0.00}$ $\mathbf{67.89\pm 0.00}$ $\mathbf{86.59\pm 0.00}$
$25$ ImageNet $11.38\pm 1.11$ $20.83\pm 2.00$ $49.29\pm 1.13$ $70.33\pm 0.29$
MoCoV3 $6.15\pm 0.00$ $30.89\pm 0.00$ $66.67\pm 0.00$ $85.37\pm 0.00$
SimCLRV2 $11.46\pm 0.00$ $33.74\pm 0.00$ $69.11\pm 0.00$ $84.15\pm 0.00$
DINO $10.32\pm 0.35$ $24.59\pm 1.02$ $55.69\pm 0.81$ $75.81\pm 0.20$
Our method $\mathbf{4.62\pm 0.00}$ $\mathbf{32.93\pm 0.00}$ $\mathbf{73.17\pm 0.00}$ $\mathbf{91.87\pm 0.00}$
$50$ ImageNet $7.36\pm 0.31$ $29.98\pm 1.73$ $67.58\pm 2.45$ $83.23\pm 0.34$
MoCoV3 $4.42\pm 0.00$ $39.43\pm 0.00$ $77.64\pm 0.00$ $92.68\pm 0.00$
SimCLRV2 $7.93\pm 0.00$ $37.40\pm 0.00$ $73.98\pm 0.00$ $91.87\pm 0.00$
DINO $7.62\pm 1.04$ $30.89\pm 1.22$ $62.81\pm 1.43$ $82.32\pm 3.05$
Our method $\mathbf{4.31\pm 0.01}$ $\mathbf{39.63\pm 0.61}$ $\mathbf{80.08\pm 0.41}$ $\mathbf{93.50\pm 0.41}$
Table 2: Comparing landmark detection test performance on Cephalometric dataset through our DDPM self-supervised pre-training and ImageNet, MoCoV3, SimCLRV2, DINO pre-training across varying numbers of labeled training images, highlighting best results in bold.
Number of Labeled Training Samples Pre-trained Models Cephalometric
MRE $\downarrow$ SDR(%) $\uparrow$
(mm) 2mm 2.5mm 3mm 4mm
$1$ ImageNet $48.69\pm 0.03$ $0.12\pm 0.01$ $0.16\pm 0.01$ $0.21\pm 0.02$ $0.41\pm 0.02$
MoCoV3 $24.39\pm 0.75$ $18.36\pm 0.35$ $23.92\pm 0.84$ $28.55\pm 0.93$ $35.89\pm 1.11$
SimCLRV2 $49.27\pm 9.00$ $4.14\pm 2.01$ $5.42\pm 2.56$ $6.73\pm 3.04$ $9.13\pm 4.10$
DINO $30.78\pm 0.73$ $7.64\pm 0.38$ $10.45\pm 0.50$ $13.57\pm 0.52$ $18.68\pm 0.64$
Our method $\mathbf{6.87\pm 0.00}$ $\mathbf{49.43\pm 0.00}$ $\mathbf{58.19\pm 0.00}$ $\mathbf{66.32\pm 0.00}$ $\mathbf{74.78\pm 0.00}$
$5$ ImageNet $46.84\pm 0.03$ $0.14\pm 0.02$ $0.22\pm 0.01$ $0.29\pm 0.00$ $0.52\pm 0.01$
MoCoV3 $7.53\pm 0.20$ $46.17\pm 0.41$ $55.26\pm 0.64$ $64.29\pm 0.53$ $75.48\pm 0.40$
SimCLRV2 $13.73\pm 0.16$ $33.69\pm 0.18$ $41.48\pm 0.35$ $49.31\pm 0.31$ $59.40\pm 0.10$
DINO $12.85\pm 0.22$ $35.88\pm 0.09$ $43.80\pm 0.27$ $50.65\pm 0.15$ $60.20\pm 0.32$
Our method $\mathbf{2.69\pm 0.00}$ $\mathbf{64.72\pm 0.00}$ $\mathbf{75.16\pm 0.00}$ $\mathbf{83.18\pm 0.00}$ $\mathbf{90.59\pm 0.00}$
$10$ ImageNet $27.44\pm 0.23$ $10.18\pm 0.03$ $13.33\pm 0.25$ $16.96\pm 0.20$ $22.68\pm 0.35$
MoCoV3 $5.32\pm 0.02$ $56.22\pm 0.29$ $66.91\pm 0.29$ $75.81\pm 0.13$ $85.25\pm 0.06$
SimCLRV2 $10.09\pm 0.07$ $43.67\pm 0.15$ $52.84\pm 0.09$ $61.19\pm 0.03$ $71.27\pm 0.20$
DINO $8.73\pm 0.08$ $48.19\pm 0.15$ $57.70\pm 0.27$ $65.90\pm 0.15$ $75.64\pm 0.00$
Our method $\mathbf{1.93\pm 0.00}$ $\mathbf{72.25\pm 0.00}$ $\mathbf{82.48\pm 0.00}$ $\mathbf{89.83\pm 0.00}$ $\mathbf{96.13\pm 0.00}$
$25$ ImageNet $4.18\pm 0.01$ $57.97\pm 0.37$ $69.03\pm 0.46$ $78.04\pm 0.20$ $87.11\pm 0.09$
MoCoV3 $2.22\pm 0.01$ $73.21\pm 0.20$ $82.69\pm 0.07$ $89.98\pm 0.25$ $95.79\pm 0.08$
SimCLRV2 $2.19\pm 0.01$ $69.35\pm 0.21$ $80.40\pm 0.04$ $88.36\pm 0.15$ $95.39\pm 0.17$
DINO $3.02\pm 0.06$ $68.22\pm 0.33$ $78.70\pm 0.27$ $86.81\pm 0.07$ $93.46\pm 0.09$
Our method $\mathbf{1.62\pm 0.03}$ $\mathbf{77.33\pm 1.12}$ $\mathbf{85.84\pm 1.49}$ $\mathbf{91.71\pm 1.20}$ $\mathbf{96.86\pm 0.63}$
$50$ ImageNet $1.60\pm 0.04$ $81.17\pm 0.42$ $89.20\pm 0.43$ $94.49\pm 0.32$ $98.29\pm 0.12$
MoCoV3 $1.63\pm 0.01$ $81.72\pm 0.04$ $89.72\pm 0.08$ $94.76\pm 0.02$ $98.28\pm 0.05$
SimCLRV2 $1.67\pm 0.02$ $79.55\pm 0.24$ $87.65\pm 0.27$ $93.67\pm 0.03$ $97.84\pm 0.01$
DINO $1.65\pm 0.01$ $80.89\pm 0.02$ $88.47\pm 0.48$ $93.75\pm 0.06$ $97.82\pm 0.07$
Our method $\mathbf{1.40\pm 0.01}$ $\mathbf{82.16\pm 1.16}$ $\mathbf{89.98\pm 0.62}$ $\mathbf{94.63\pm 0.54}$ $\mathbf{98.30\pm 0.09}$
Table 3: Comparing landmark detection test performance on the Hand dataset through our DDPM self-supervised pre-training and ImageNet, MoCoV3, SimCLRV2, and DINO pre-training across varying numbers of labeled training images, highlighting best results in bold.
Number of Labeled Training Samples Pre-trained Models Hand
MRE $\downarrow$ SDR(%) $\uparrow$
(mm) 2mm 4mm 10mm
$1$ ImageNet $79.32\pm 0.03$ $0.04\pm 0.00$ $0.14\pm 0.00$ $1.08\pm 0.02$
MoCoV3 $100.17\pm 0.10$ $0.42\pm 0.02$ $0.83\pm 0.02$ $1.60\pm 0.03$
SimCLRV2 $98.52\pm 0.09$ $0.06\pm 0.01$ $0.31\pm 0.00$ $1.02\pm 0.04$
DINO $97.31\pm 0.09$ $0.47\pm 0.04$ $1.27\pm 0.03$ $3.30\pm 0.18$
Our method $\mathbf{28.75\pm 0.00}$ $\mathbf{27.87\pm 0.00}$ $\mathbf{46.44\pm 0.00}$ $\mathbf{58.75\pm 0.00}$
$5$ ImageNet $86.12\pm 0.03$ $0.01\pm 0.01$ $0.09\pm 0.01$ $1.01\pm 0.01$
MoCoV3 $61.20\pm 0.39$ $9.82\pm 0.16$ $18.94\pm 0.29$ $32.11\pm 0.61$
SimCLRV2 $75.07\pm 0.48$ $5.99\pm 0.07$ $11.87\pm 0.14$ $20.35\pm 0.34$
DINO $86.03\pm 0.64$ $2.05\pm 0.31$ $5.79\pm 0.55$ $11.96\pm 0.84$
Our method $\mathbf{6.86\pm 0.00}$ $\mathbf{61.09\pm 0.00}$ $\mathbf{83.30\pm 0.00}$ $\mathbf{90.62\pm 0.00}$
$10$ ImageNet $85.87\pm 0.00$ $0.01\pm 0.00$ $0.15\pm 0.00$ $0.99\pm 0.00$
MoCoV3 $55.43\pm 0.29$ $13.92\pm 0.17$ $26.00\pm 0.15$ $40.21\pm 0.20$
SimCLRV2 $55.58\pm 0.75$ $16.72\pm 0.36$ $29.45\pm 0.45$ $42.18\pm 0.72$
DINO $60.58\pm 0.53$ $11.09\pm 0.14$ $23.00\pm 0.12$ $35.47\pm 0.32$
Our method $\mathbf{4.94\pm 0.11}$ $\mathbf{62.55\pm 0.55}$ $\mathbf{86.08\pm 0.25}$ $\mathbf{94.28\pm 0.17}$
$25$ ImageNet $24.96\pm 0.47$ $42.79\pm 0.60$ $57.93\pm 0.28$ $71.68\pm 0.18$
MoCoV3 $40.03\pm 0.12$ $27.19\pm 0.09$ $44.30\pm 0.27$ $56.55\pm 0.05$
SimCLRV2 $40.02\pm 0.11$ $29.69\pm 0.06$ $47.03\pm 0.00$ $59.73\pm 0.11$
DINO $49.27\pm 0.30$ $21.09\pm 0.06$ $36.12\pm 0.10$ $48.76\pm 0.20$
Our method $\mathbf{2.39\pm 0.00}$ $\mathbf{72.99\pm 0.00}$ $\mathbf{93.18\pm 0.00}$ $\mathbf{97.87\pm 0.00}$
$50$ ImageNet $10.79\pm 0.23$ $61.98\pm 0.36$ $78.51\pm 0.26$ $90.04\pm 0.28$
MoCoV3 $23.77\pm 0.34$ $47.44\pm 0.10$ $67.96\pm 0.14$ $76.18\pm 0.23$
SimCLRV2 $23.95\pm 0.18$ $49.28\pm 0.30$ $68.56\pm 0.45$ $77.03\pm 0.69$
DINO $34.95\pm 0.59$ $34.06\pm 0.97$ $52.32\pm 1.20$ $64.21\pm 0.80$
Our method $\mathbf{1.74\pm 0.03}$ $\mathbf{78.04\pm 0.94}$ $\mathbf{96.35\pm 0.51}$ $\mathbf{99.02\pm 0.04}$
Figure 4: Comparison of landmark detection performance between our pre-training method and other self-supervised and ImageNet pre-training approaches on Chest, Cephalometric, and Hand x-ray test sets, using only 10 labeled training samples.
Table 4: Landmark detection test performance of our method, pre-trained on the Hand dataset and fine-tuned on the Chest / Cephalometric datasets, across varying numbers of labeled training images.
Number of Labeled Training Samples Hand pre-training $\rightarrow$ Chest fine-tuning Hand pre-training $\rightarrow$ Cephalometric fine-tuning
MRE±$\pm$±STD $\downarrow$ SDR(%)$\pm$±STD $\uparrow$ MRE±$\pm$±STD $\downarrow$ SDR(%)$\pm$±STD $\uparrow$
(px) 3px 6px 9px (mm) 2mm 2.5mm 3mm 4mm
$1$ $23.79\pm 0.00$ $14.23\pm 0.00$ $35.37\pm 0.00$ $52.03\pm 0.00$ $9.56\pm 0.28$ $30.45\pm 1.73$ $38.60\pm 1.78$ $47.75\pm 2.00$ $60.21\pm 1.73$
$5$ $5.86\pm 0.00$ $32.93\pm 0.00$ $65.04\pm 0.00$ $82.93\pm 0.00$ $5.42\pm 0.00$ $55.07\pm 0.00$ $64.55\pm 0.00$ $72.44\pm 0.00$ $79.71\pm 0.00$
$10$ $5.08\pm 0.00$ $34.55\pm 0.00$ $70.73\pm 0.00$ $85.37\pm 0.00$ $2.78\pm 0.00$ $63.52\pm 0.30$ $74.20\pm 0.62$ $82.50\pm 0.35$ $89.97\pm 0.35$
$25$ $4.39\pm 0.00$ $38.62\pm 0.00$ $77.24\pm 0.00$ $91.46\pm 0.00$ $1.70\pm 0.00$ $74.03\pm 0.01$ $83.97\pm 0.05$ $90.95\pm 0.04$ $96.26\pm 0.03$
$50$ $3.99\pm 0.00$ $41.87\pm 0.00$ $82.52\pm 0.00$ $92.68\pm 0.00$ $1.50\pm 0.01$ $79.90\pm 0.02$ $87.85\pm 0.02$ $93.12\pm 0.00$ $97.37\pm 0.00$
5.3 Impact of different pre-training datasets
In this section, we explore whether a different in-domain dataset can still boost the performance in the potential situation where both the number of available images and annotations are limited. To this purpose, we select the best DDPM model pre-trained on the larger dataset (Hand), and we fine-tune it on the other two smaller datasets (Chest and Cephalometric), with the same protocol as the previous experiments (three runs and different numbers of annotated training images for the downstream task). Table 4 reports the obtained results. Our experiments reveal the efficacy of using a larger in-domain dataset (Hand) for pre-training in scenarios with limited target task data.
Using a model pre-trained on the Hand dataset for the Chest landmark detection task yields better results compared to a model pre-trained on the Chest dataset itself for most training samples quantity scenarios (5 to 50 labeled images). Even with just one labeled sample, the Hand pre-trained model achieves comparable performance to the Chest pre-trained model.
For the Cephalometric landmark detection task, the Hand pre-trained model does not outperform the Cephalometric pre-trained model, but it does achieve competitive results across all the selected number of training samples. The performance gap between the two pre-training approaches is relatively small, indicating that the Hand pre-training still transfers useful features to the Cephalometric task. Importantly, our approach consistently outperforms other pre-training methods (ImageNet, MoCoV3, SimCLRV2, DINO), regardless of the number of labeled training samples, for both tasks. This underscores the value of in-domain pre-training using our approach, even when the pre-training dataset differs from the target task. Our goal is to utilize images from the same dataset as the downstream task, thereby simulating real-world scenarios with limited data and label availability. Furthermore, we focused on label-efficient pre-training while minimizing computational resource usage, given that training diffusion models on extensive unlabeled datasets demands substantial resources and time. Even though different larger unlabeled datasets might enhance pre-training outcomes, our main point is that using a limited unannotated dataset can still surpass self-supervised and supervised state-of-the-art pre-training techniques like ImageNet (which has one million images), when images and annotations availability are constrained.
6 Conclusions
Despite the widespread application of deep neural networks for medical image analysis, data scarcity in images and annotations represents a fundamental challenge for this domain. In this research, we propose and evaluate the impact of a pre-training approach based on self-supervision with DDPM for label-efficient landmark detection tasks in x-ray images, benchmarking it against the typically adopted supervised ImageNet pre-training and self-supervised MoCoV3, SimCLRV2, and DINO. Specifically, given an x-ray image dataset, we leverage the training set in a self-supervised fashion for pre-training a DDPM based on a Unet, further fine-tuning it on the available annotated training images. Our results on three popular benchmark datasets show that the proposed method outperforms ImageNet supervised pre-training and MoCoV3, SimCLRV2, DINO self-supervised pre-training in the context of few annotations available for the landmark detection task. Moreover, such improvement seems to be preserved when pre-training on a different in-domain dataset and fine-tuning on the downstream one, simulating the interesting case where both annotations and images are limited for the downstream task. We believe that the proposed method showcases the potential of self-supervised diffusion-based pre-training for the landmark detection task, potentially opening new research directions to effectively address data scarcity challenges in this domain.